OpenStudy (anonymous):

A heat engine operates between 500C and 250C. a) Calculate its maximum effeciency. (ans: 32.3%) b) How much heat must be input to produce a work output of 750 J ? (ans: 2320 J)

OpenStudy (anonymous):

a) see Carnot cycle \[\eta=1-T_c/T_h=50\]

OpenStudy (anonymous):

that's is what i thought, but the answer provided says answer is 32.3% Also, what about part b?

OpenStudy (anonymous):

\[\eta=W/Q\]

OpenStudy (anonymous):

using given 23% you will get that Q. but I think 50% is right

OpenStudy (anonymous):

sorry 32.2%

OpenStudy (anonymous):

with the information given it has to be 50? must be a mistake

OpenStudy (anonymous):

right

OpenStudy (anonymous):

hey man thanks I really appreciate your time. just needed some validation.

OpenStudy (anonymous):

you are welcome